OpenStudy (anonymous):

help?!?!(medal will be given) Calculate the area of triangle QRS with altitude ST, given Q (0, 5), R (−5, 0), S (−3, 4), and T (−2, 3). 6.2 square units 7 square units 5.9 square units 5 square units

OpenStudy (anonymous):

@esshotwired @phi help plz

OpenStudy (anonymous):

@esshotwired

OpenStudy (anonymous):

@phi

OpenStudy (phi):

Did you plot the points to see what it looks like ?

OpenStudy (anonymous):

it looks like triangle

OpenStudy (phi):

Here is a graph of the problem to find the area of the triangle, you need to find the length of its base QR use the distance formula and the length of its altitude ST use the distance formula can you do that ?

OpenStudy (anonymous):

QR= (0,5),(-5,0) =SQUAR ROOT (X2-X1)2+(Y2-Y1)2 (-5-0)2+(0-5)2=25+25=50=7.071

OpenStudy (phi):

yes, but in this case, you might want to write it as \[ \sqrt{50} = \sqrt{2\cdot 25} = 5\sqrt{2} \] do the same for ST. (leave it in radical form) then do area (of a triangle) = 1/2 * base * height

OpenStudy (anonymous):

ST=(-3,4),(-2,3) -2--3^2+3-4^2= 1+1=2

OpenStudy (phi):

don't forget the square root

OpenStudy (anonymous):

SQUAR ROOT 2 =1.41

OpenStudy (phi):

but leave it \( \sqrt{2} \) now what is the area? \[ A= \frac{1}{2} \cdot base \cdot { height } \\ = \frac{1}{2} \cdot 5 \sqrt{2} \cdot \sqrt{2} \] notice that \( \sqrt{2} \cdot \sqrt{2} = 2\)

OpenStudy (anonymous):

5

OpenStudy (anonymous):

THX FOR THE HELP

OpenStudy (phi):

yw